# Uncertainty Visualization via Low-Dimensional Posterior Projections

### v1

@@ B.3 Projected posterior distribution
Appendix C PPD learning with an EBM
@@
$\tilde{\mathbf{x}}=\alpha_{\mathbf{t}}\mathbf{x}+\sqrt{1-\alpha_{\mathbf{t}}^{2}}\,\mathbf{n}.$ (30)
@@
$f_{t}(\tilde{\bm{x}};\bm{\theta})\approx\log p_{\tilde{\mathbf{x}},\textnormal{t}}(\tilde{\bm{x}},t)=\log p_{\tilde{\mathbf{x}}|\textnormal{t}}(\tilde{\bm{x}}|t)+\log p_{\textnormal{t}}(t).$ (31)
@@
$f_{1}(\bm{x};\bm{\theta})\approx\log p_{\mathbf{x}}(\bm{x})+\log p_{\textnormal{t}}(1).$ (32)
@@
$\mathrm{Softmax}(\bm{f}(\tilde{\bm{x}};\bm{\theta}))_{t}\approx\frac{p_{\tilde{\mathbf{x}},\textnormal{t}}(\tilde{\bm{x}},t)}{\sum_{t^{\prime}=1}^{T}p_{\tilde{\mathbf{x}},\textnormal{t}}(\tilde{\bm{x}},t^{\prime})}=p_{\textnormal{t}|\tilde{\mathbf{x}}}(t|\tilde{\bm{x}}).$ (33)
@@
$\qquad\left.+\mathrm{Softmax}(\bm{f}(\tilde{\bm{x}};\bm{\theta}))_{t}\right],$ (34)
@@
Appendix D Datasets and restoration tasks
We evaluated our method on a range of datasets and tasks spanning common scenarios in low-level vision. Specifically, we used the MNIST dataset [13] ($28\times 28$), the CelebA dataset [36] (cropped and resized to $160\times 160$ as in SRFlow [38]), the CelebA-HQ dataset [28] ($256\times 256$), and the ImageNet 1K dataset [57] ($128\times 128$). The tasks we experimented with were:
@@ Appendix D Datasets and restoration tasks
Appendix E Architectures and training hyperparameters
@@
All networks were trained using the Adam optimizer, stopping the training process once it reached a minimal value of the objective over a dedicated validation set. For the training of the PC predictor and the PPD model, the learning rate was kept fixed. On the other hand, for training the MMSE estimator, the learning rate was reduced by half every 5000 steps. Table 2 summarizes the architectures and training hyperparameters used for the networks in each task/dataset.
Table 2: Architectures, learning rates, and number of steps used per task/dataset.
@@
Appendix F Additional Results
F.1 PPDs
Figures 11-16 provide more examples of PPDs captured by PPDE for the datasets and tasks presented in the main text.
Figure 11: More examples of MNIST inpainting PPDs.
Figure 12: More examples of MNIST denoising.
Figure 13: More examples of CelebA-HQ inpainting.
Figure 14: More examples of CelebA-HQ colorization.
Figure 15: More examples of CelebA 8$\times$ super-rsolution.
Figure 16: More examples of ImageNet colorizaion.
F.2 Comparisons to NPPC and KDE
Figure 17: More comparisons of MNIST inpainting PPDs. The posterior samples on the right were obtained using an EBM trained on MNIST.
Figure 18: More comparisons of MNIST denoising. The posterior samples on the right were obtained using an EBM trained on MNIST.
Figure 19: More comparisons of CelebA-HQ inpainting. The posterior samples on the right were obtained using MAT [33].
Figure 20: More comparisons of CelebA-HQ colorization. The posterior samples on the right were obtained using DDNM [72].
Figure 21: More comparisons of CelebA 8$\times$ super-resolution.
Figure 22: More comparisons of ImageNet colorization.
Figures 17-22 provide further negative log-likelihood comparisons of PPDE with the respective baselines.

### v2

@@ A.3 Projected posterior distribution
Appendix B PPD learning with an EBM
@@
$\tilde{\mathbf{x}}=\alpha_{\mathbf{t}}\mathbf{x}+\sqrt{1-\alpha_{\mathbf{t}}^{2}}\,\mathbf{n}.$ (S21)
@@
$f_{t}(\tilde{\bm{x}};\bm{\theta})\approx\log p_{\tilde{\mathbf{x}},\textnormal{t}}(\tilde{\bm{x}},t)=\log p_{\tilde{\mathbf{x}}|\textnormal{t}}(\tilde{\bm{x}}|t)+\log p_{\textnormal{t}}(t).$ (S22)
@@
$f_{1}(\bm{x};\bm{\theta})\approx\log p_{\mathbf{x}}(\bm{x})+\log p_{\textnormal{t}}(1).$ (S23)
@@
$\mathrm{Softmax}(\bm{f}(\tilde{\bm{x}};\bm{\theta}))_{t}\approx\frac{p_{\tilde{\mathbf{x}},\textnormal{t}}(\tilde{\bm{x}},t)}{\sum_{t^{\prime}=1}^{T}p_{\tilde{\mathbf{x}},\textnormal{t}}(\tilde{\bm{x}},t^{\prime})}=p_{\textnormal{t}|\tilde{\mathbf{x}}}(t|\tilde{\bm{x}}).$ (S24)
@@
$\qquad\left.+\mathrm{Softmax}(\bm{f}(\tilde{\bm{x}};\bm{\theta}))_{t}\right],$ (S25)
@@
Appendix C Datasets and restoration tasks
We evaluated our method on a range of datasets and tasks spanning common scenarios in low-level vision. Specifically, we used the MNIST dataset [13] ($28\times 28$), the CelebA dataset [35] (cropped and resized to $160\times 160$ as in SRFlow [37]), the CelebA-HQ dataset [28] ($256\times 256$), the ImageNet 1K dataset [56] ($128\times 128$) and an a cells microscopy dataset presented in [70]. The tasks we experimented with were:
@@ Appendix C Datasets and restoration tasks
  • Biological image-to-image. Transforming a given microscopic biological imaged specimen with one fluorescent dye appears as if it was imaged by another. We have broken this task into transforming patches of size $64\times 64$ between the 2 domains.
Appendix D Architectures and training hyperparameters
@@
All networks were trained using the Adam optimizer, stopping the training process once it reached a minimal value of the objective over a dedicated validation set. For the training of the PC predictor and the PPD model, the learning rate was kept fixed. On the other hand, for training the MMSE estimator, the learning rate was reduced by half every 5000 steps. Table S2 summarizes the architectures and training hyperparameters used for the networks in each task/dataset.
Table S2: Architectures, learning rates, and number of steps used per task/dataset.
@@
Biological
Image-to-image Pretrained U-Net-M $10^{-4}$ 81,100 cEBM-L $10^{-4}$ 402,800
Appendix E Additional Results
E.1 PPDs
Figures S12-S18 provide more examples of PPDs captured by PPDE for the datasets and tasks presented in the main text.
Figure S12: More examples of MNIST inpainting PPDs.
Figure S13: More examples of MNIST denoising.
Figure S14: More examples of CelebA-HQ colorization.
Figure S15: More examples of CelebA-HQ inpainting.
Figure S16: More examples of CelebA 8$\times$ super-resolution.
Figure S17: More examples of ImageNet colorizaion.
Figure S18: More examples of Biological image-to-image transfer.
E.2 Comparisons to NPPC and KDE
Figure S19: More comparisons of MNIST inpainting PPDs. The posterior samples on the right were obtained using an EBM trained on MNIST.
Figure S20: More comparisons of MNIST denoising. The posterior samples on the right were obtained using an EBM trained on MNIST.
Figure S21: More comparisons of CelebA-HQ inpainting. The posterior samples on the right were obtained using DPS [33].
Figure S22: More comparisons of CelebA-HQ colorization. The posterior samples on the right were obtained using DDNM [72].
Figure S23: More comparisons of CelebA 8$\times$ super-resolution.
Figure S24: More comparisons of ImageNet colorization.
Figures S19-S24 provide further negative log-likelihood comparisons of PPDE with the respective baselines.